Add the following vectors. Find the size (magnitude) of the resultant vector. Determine the angle from the resultant vector counterclockwise to the right side x-axis.

Add a vector 8 units due south and a vector 6 units due east.

Show your answer on a graph.
Answer #1

both vectors are perpendicular to each other so

magnitude = sqrt(6^2 + 8^2) = 10 units

direction = tan^-1 ( 8/6) = 53.13o

direction counterclockwise of x-axis = 360 - 53.13 = 306.87o